OpenStudy (anonymous):

Write an equation of the line containing the given point and parallel to the given line. (5,-6); 4x-5y=9 the equation of the line is y=

OpenStudy (anonymous):

so first put it into slope intercept form which is y=mx+b

OpenStudy (anonymous):

i dont know how to do that :/

OpenStudy (anonymous):

took me a while too :) so you want to isolate the y

OpenStudy (anonymous):

start by subtracting the 4x from both sides this cancels it out on the left hand of the = sign and puts it on the right with the 9 like so -5y=9-4x

OpenStudy (anonymous):

you can then rearrange to make it look more like the final result giving yourself -5y=-4x+9 this works for aesthetics but if the (-) sign gets overwhelming you can wait until later too

OpenStudy (anonymous):

kk thank you, and i think you divide both sides by -5

OpenStudy (anonymous):

Yup then divide both by -5 and you can keep it as y=\[\frac{ 4x-9 }{ 5 }\] or make it \[y=\frac{ 4 }{ 5}x-\frac{ 9 }{ 5}\] whichever is easier for you to understand I personally understand the 2nd one better

OpenStudy (anonymous):

The second part is finding a line parallel which uses the same slope in this case \[\frac{ 4 }{ 5 }\]

OpenStudy (anonymous):

to have it go through the point given you just put those into your equation of \[y=\frac{ 4 }{ 5}x+b\] so x=5 and y=-6 this gives you \[-6=\frac{ 4 }{ 5}*5+b\] solve for b

OpenStudy (anonymous):

kk thank you!! i really apreciate it :)

OpenStudy (anonymous):

b=-10 so final answer should be \[y=\frac{ 4 }{ 5 }x-10\] no problem! Glad to help

OpenStudy (anonymous):

do you need help with a perpendicular one too?

OpenStudy (anonymous):

no just with the equations for now :)